(3,-4),y=5x-3
solve for parallel line

Answer by mananth(16946)

y = 5 x -3
Compare this equation with y=mx+b
slope m = 5 0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be 5
m= 5 ,point ( 3 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4 = 15.00 + b
b= -19
m= 5
Plug value of the slope and b in y = mx +b
The required equation is y = 5 x -19
y = 5 x -19
m.ananth@hotmail.ca